एक पिण्ड प्रथम 2 सैकेंड में 200मीटर की दूरी तय करती है अगले 4सैकेन्ड में 220मीटर की दूरी तय करती है तो प्रारम्भ से 7वें सैकेंड के अंत में पिण्ड का वेग करता होगा

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Using s = ut + 0.5 a t^2

Equation 1 is

200 = 2u + 0.5( a ) ( 4)

And equation 2 is

S6-s4 = ( u6-0.5a*36) - ( u2+ 0.5a*4)

Simplify and use these equations to obtain u and a and then find the required data

Regards
MAHIPAL VISHNOI
one year ago
Make the equation according to question follow...
S =ut+1/2 at2
 
1.   S=200m ,t=2s
Then 200=2u+1/2 a(4)
     100 =  u+a ...............(1)
2. S=(200+220)=420m,. t=(2+4)=6s
So  420=6u+1/2a(36)
       70=u+3a................(2)
 
By equation 1&2
a= —15.    u=115
 
Now,    V=?    For  t=7
By    V= U+at
         V=115+(15)7
        V=10m/s
 
Solved By🎤📝     MAHIPAL  VISHNOI 
                                 +919511361229
            Instagram:- mahi.official1229